# On-device Collaborative Language Modeling via a Mixture of Generalists and Specialists

### v1

@@ Abstract
We target on-device collaborative fine-tuning of Large Language Models (LLMs) by adapting a Mixture of Experts (MoE) architecture, where experts are Low-Rank Adaptation (LoRA) modules. In conventional MoE approaches, experts develop into specialists throughout training. In contrast, we propose a novel Collaborative learning approach via a Mixture of Generalists and Specialists (CoMiGS). Diversifying into the two roles is achieved by aggregating certain experts globally while keeping others localized to specialize in user-specific datasets. Central to our work is a learnable routing network that routes at a token level, balancing collaboration and personalization at the finest granularity. Our method consistently demonstrates superior performance in scenarios with high data heterogeneity across various datasets. By design, our approach accommodates varying computational resource constraints among users as shown in different numbers of LoRA experts. We further showcase that low-resourced users can benefit from high-resourced users with high data quantity.
@@ 1 Introduction
Large Language Models (LLMs) have been showing great success serving as foundation models, evidenced by their capability to understand a wide range of tasks. Their applicability and personalized adaptation to end users can further be enhanced by fine-tuning on device. On-device fine-tuning of LLMs poses new challenges such as restricted and varying computational resources, scarce and heterogenous local data, and privacy concerns related to data sharing. As a solution, we study personalized collaborative on-device fine-tuning of LLMs in this work.
A parameter-efficient method for performing LLM fine-tuning is through Low-Rank Adaptors (LoRA, Hu et al. (2021)), where model updates ($\Delta{\bm{W}}\in\mathbb{R}^{m\times n})$ are approximated by the multiplication of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{m\times r}$ and ${\bm{B}}\in\mathbb{R}^{r\times n}(r\ll m,n)$. However, when system heterogeneity arises—meaning users are equipped with different-sized LoRA modules—collaboration through simple model parameter aggregation becomes unfeasible. Previous works have proposed various techniques for aggregating LoRA modules of different ranks (Cho et al., 2023; Bai et al., 2024). In this work, we address system heterogeneity by allowing users to have varying numbers of LoRA modules. That is, high-resourced users are equipped with a larger number of LoRA modules locally.
The aggregation of LoRA modules within a user can be done through a Mixture-of-Experts (MoE) fashion, where each LoRA module is regarded as an expert, as in Wu et al. (2024); Zadouri et al. (2024). Typically, a Multi-layer Perceptron (MLP) routing network dynamically assigns input units to each expert, and the expert scores serve as aggregation weights to combine expert outputs. Routing is often done at the token level, a design choice proven to bring optimal performances (Jiang et al., 2024; Fan et al., 2024b). Tokens within an input sequence are directed to various experts based on the routing mechanism. Since the routing and expert networks are jointly trained, the diversification of experts occurs naturally. It is observed that different experts develop distinct context-independent specializations in syntax and locality. (Jiang et al., 2024; Xue et al., 2024). We call these experts specialists. While specialist experts show promising results, can the performances be further advanced by introducing some experts with more general knowledge (generalists)?
The modular nature of MoEs facilitates the coexistence of generalists and specialists in a collaborative learning setup. It allows for the separation of certain experts, whose parameters can be aggregated across users (generalists), while other experts remain localized to specialize in user-specific data (specialists). Presumably, generalists are of importance, particularly in language modeling. End users, such as mobile device users, may have unique topics or language patterns but frequently share common tokens in their vocabularies (see Figure 5). Jointly learning these shared tokens should be more efficient than each user learning them alone.
Additionally, MoE architecture can deal with system heterogeneity nicely, allowing users to have varying numbers of specialists as long as the number of generalist experts remains consistent for aggregation. This enables high-resourced users to leverage more experts for their local data, while low-resourced users still benefit from shared expertise.
Balancing personalization and collaboration is a challenging task in collaborative learning. Typically, the challenge is addressed at a client level, where a pairwise collaboration graph is inferred from gradient or prediction similarity across users (Fan et al., 2024a; Wagner et al., 2024). We take a novel approach by addressing it at a token level. Our approach leverages the powerful routing mechanism of MoEs, where the routing results are used to weigh generalist and specialist experts for each input token. The finest-grained approach should offer the most tailored solution to the end users.
@@
  • We propose a novel approach (CoMiGS) for on-device personalized collaborative fine-tuning of LLMs, by introducing a mixture of generalist and specialist LoRA experts. The approach balances personalization and collaboration at the finest granularity.
@@
  • Our framework separates resource heterogeneity from data quantity. Users with larger datasets benefit from a bigger model, while high-resourced users with smaller datasets are less prone to overfitting.
@@ 2 Related Work
Figure 1: Diagram of our proposed method illustrated with a 4-client setup, with a dark blue frame denoting a block. Within each block, generalist experts ($\{\theta^{G}_{1},{\mathbf{\theta}}^{G}_{2},{\mathbf{\theta}}^{G}_{3},{\mathbf{\theta}}^{G}_{4}\}$) are aggregated across users, and specialist experts ($\{{\mathbf{\theta}}^{Si}_{1},{\mathbf{\theta}}^{Si}_{2},{\mathbf{\theta}}^{Si}_{3},{\mathbf{\theta}}^{Si}_{4}\}$) are kept local. Router and expert parameters are updated iteratively using local validation and training datasets.
@@ 3 Method
Aiming to achieve better personalized performance for each user, we propose leveraging collective knowledge while maintaining specialized expertise within each user through a mixture of generalist and specialist experts. As a solution, we introduce a multi-round, gradient-based algorithm based on the following key insights: (1) generalist models emerge through the aggregation of weights during communication rounds, and (2) expert and routing parameters are updated alternately, guided by two distinct and disentangled loss functions. We formalize our problem setup and solution as follows:
@@ 3.1 Notions and problem setup
As illustrated in Figure 1, there are two sets of model parameters within each user: expert parameters, denoted as ${\mathbf{\Theta}}={\mathbf{\theta}}^{G}\cup\{{\mathbf{\theta}}_{i}^{S}\}$, where ${\mathbf{\theta}}^{G}$ is shared across the users and $\{{\mathbf{\theta}}_{i}^{S}\}$ are user-specific specialist parameters; and routing parameters, denoted as ${\mathbf{\Phi}}=\{{\mathbf{\phi}}_{i}\}$. $i\in[0,N)$ is the user index. The ${\mathbf{\theta}}$s are a union of LoRA A and B matrices. Our ultimate goal is to optimize the average performance of users, following the averaged personalized learning objectives on user-specific targets:
$\min_{{\mathbf{\Theta}},{\mathbf{\Phi}}}\frac{1}{N}\sum_{i}{\mathcal{L}}(f({\bm{X}}_{i}^{\text{test}};{\mathbf{\theta}}^{G},{\mathbf{\theta}}^{S}_{i},{\mathbf{\phi}}_{i}),{\bm{X}}_{i}^{\text{test}})$ (1)
where ${\mathcal{L}}$ is the language modeling loss. Note we write ${\bm{X}}_{i}^{\text{test}}$ as the label here, as this is a self-supervised task. Labels are simply shifted inputs.
@@
Input: Expert parameters $\{{\mathbf{\theta}}_{i,0}^{G},{\mathbf{\theta}}_{i,0}^{S}\}$, routing parameters ${\mathbf{\phi}}_{i,0}$}. Local training data and validation data $\{{\bm{X}}_{i}^{\text{train}},{\bm{X}}_{i}^{\text{valid}}\}$, $i\in[0,N)$. Communication round $T$ and routing update period $\tau$. Load balancing weight $\lambda$.
@@
     Each device $i\in[0,N)$ sends generalist weights ${\mathbf{\theta}}_{i,t}^{G}$ to the server
@@
3.2 Our approach
Since the objective in (1) is unattainable due to the unavailability of test loss during training, we approximate it using validation loss. We further reformulate our proxy objective as a bi-level optimization problem, as presented in (4). The routing parameters ${\mathbf{\Phi}}=\{{\mathbf{\phi}}_{i}\}$ are updated based on the validation loss, which reflects the target distribution (outer optimization), while the expert parameters ${\mathbf{\Theta}}={\mathbf{\theta}}^{G}\cup\{{\mathbf{\theta}}^{S}_{i}\}$ are updated using the training loss (inner optimization). This formulation arises from the idea that expert knowledge can be effectively developed from the large training sets, while how to combine that knowledge should remain target-specific.
To solve (4), we use a multi-round gradient-based algorithm as shown in Alg.1, where only generalist parameters are shared and aggregated, and specialist parameters and router weights are updated locally. While the scheme requires a server, it can alternatively be implemented in a serverless all2all fashion, which requires $N$ times more communication overhead and we do not further pursue this here.
3.3 Algorithmic highlights
@@
Alternating update of two sets of parameters is a standard way to solve bi-level optimization problems. In between two communication rounds, we perform alternating updates of expert and routing parameters using local training and validation sets separately. The updates optimize the objectives given in (2) and (3) respectively. Since the updates of ${\mathbf{\Theta}}$ and ${\mathbf{\Phi}}$ are disentangled, they do not need to be updated at the same frequency. The routing parameters are smaller in size and thus can be updated less frequently. An empirical evidence is provided in Fig. 9.
When updating expert parameters, we include an additional load-balancing term as in Fedus et al. (2022), which is standard in MoE implementation and encourages even distribution of token assignments to experts. A discussion over the load balancing term is included in Appendix C.3. It is observed that a load-balancing term can improve test performance compared to not having one. However, directing more tokens to the generalists has no noticeable effect.
@@ 4.1.1 Datasets
We selected three diverse distributed datasets to showcase the efficacy of our proposed algorithm:
@@
    SlimPajama: we pick the following four categories – StackExchange, Github, ArXiv, Book from Soboleva et al. (2023);
@@
Opting for the most challenging scenario, each user is assigned a unique category, as shown in Figure 2, where users can have varying data quantities. Given our emphasis on next token prediction, we anticipate shared predictions among users while maintaining category-specific distinctions. For details of our user data splits, please refer to Appendix B.
@@ 4.1.2 Experimental Details
We start by equipping users with the same model architecture locally, to study the powerfulness of our routing mechanism. In our setup every user has two LoRA experts locally. We compare our one generalist one specialist (CoMiGS-1G1S) method to the following baselines. In order to match the trainable parameter count of our method, we use 2 times LoRA modules within each user.
@@
    • CoMiGS-2S: Both of the LoRA experts are specialists, that is, their weights are never shared and aggregated. The routing parameters are updated using a separate validation set like CoMiGS-1G1S.
    • CoMiGS-2G: Both of the LoRA experts are generalists, that is, their weights are always shared and aggregated. The routing parameters are updated using a separate validation set like CoMiGS-1G1S.
The comparison between ii) and iii) serves to illustrate the role of our routing mechanism, while the comparison between our CoMiGS-1G1S and iii) demonstrates the benefit of the coexistence of both specialists and generalists. To show the wide applicability of our method, we focus on the following two tasks:
In-Domain Tasks.
Within each user, we create validation/test datasets from the same domain as the training sets. Here we address the following two scenarios: i) diversity in language usage across users (Multilingual Wikipedia) ; ii) diversity in topic coverage across users (SlimPajama).
Out-of-Domain Tasks.
Within each user, we create validation/test datasets from a different distribution than in the training sets. During training, each user has one News category from AG News. However, the validation and test sets for each user are a homogeneous mixture of all categories. This is to account for possible topic changes within users.

### v2

@@ Abstract
On-device LLMs have gained increasing attention for their ability to enhance privacy and provide a personalized user experience. To facilitate learning with private and scarce local data, federated learning has become a standard approach, though it introduces challenges related to system and data heterogeneity among end users. As a solution, we propose a novel Collaborative learning approach with a Mixture of Generalists and Specialists (CoMiGS), being the first to effectively address both. Our approach distinguishes generalists and specialists by aggregating certain experts across end users while keeping others localized to specialize in user-specific datasets. A key innovation of our method is the bi-level optimization formulation of the Mixture-of-Experts learning objective, where the router is updated using a separate validation set that represents the target distribution. CoMiGS effectively balances collaboration and personalization, as demonstrated by its superior performance in scenarios with high data heterogeneity across multiple datasets. By design, our approach accommodates users’ varying computational resources through different numbers of specialists. By decoupling resource abundance from data quantity, CoMiGS remains robust against overfitting—due to the generalists’ regularizing effect—while adapting to local data through specialist expertise.
@@ 1 Introduction
Large Language Models (LLMs) have been showing great success serving as foundation models, evidenced by their capability to understand a wide range of tasks, such as ChatGPT (OpenAI, 2023), Claude (Anthropic, 2023), Gemini (DeepMind, 2023) and etc. However, cloud-based inference introduces significant delays for end users, and it often fails to meet their personalized needs (Ding et al., 2024; Iyengar & Adusumilli, 2024). Recently, there has been growing interest in deploying LLMs on edge devices, which offer benefits like lower latency, data localization, and more personalized user experiences (Xu et al., 2024). For instance, Apple (2024) recently launched on-device foundation models as part of its personal intelligence system.
On-device LLMs present challenges such as limited and variable computational resources, scarce and heterogeneous local data, and privacy concerns related to data sharing (Peng et al., 2024; Wagner et al., 2024). Fine-tuning is typically performed on-device to quickly adapt to users’ individual needs. While data sharing is a common solution to address local data scarcity, on-device data is often privacy-sensitive and must remain on the device. To overcome this, federated learning has been proposed as a method for enabling collaborative learning while preserving user privacy, allowing end users to collaborate by sharing model parameters (Chen et al., 2023; Zhang et al., 2023).
Federated fine-tuning of LLMs is predominately done through Low-Rank Adaptation (LoRA, Hu et al. (2021)) due to its lightweight nature so that the communication costs can be largely mitigated. Yet end devices may have different capacities, resulting in different LoRA ranks or different numbers of LoRA modules allowed on devices. Previous works have proposed various techniques for aggregating LoRA modules of different ranks (Cho et al., 2023; Bai et al., 2024). However, in both works, the devices are only equipped with shared knowledge, which makes the methods unsuitable when there is data heterogeneity across users. In such cases, a more personalized solution is needed.
End users’ local data distributions can exhibit significant statistical heterogeneity. For instance, mobile device users may have distinct linguistic habits, topic preferences, or language usage patterns, leading to widely varying word distributions. As a result, personalized solutions are necessary. Wagner et al. (2024) explored three personalized collaborator selection protocols, allowing each end user to choose their collaborators. Although these protocols effectively address data heterogeneity, they depend on model aggregation, which can only occur when users share the same model architecture.
Figure 1: Diagram of our proposed method CoMiGS illustrated with a 4-client setup, with a dark blue frame denoting a block. Within each block, generalist experts ($\{\theta^{G}_{1},{\mathbf{\theta}}^{G}_{2},{\mathbf{\theta}}^{G}_{3},{\mathbf{\theta}}^{G}_{4}\}$) are aggregated across users, and specialist experts ($\{{\mathbf{\theta}}^{Si}_{1},{\mathbf{\theta}}^{Si}_{2},{\mathbf{\theta}}^{Si}_{3},{\mathbf{\theta}}^{Si}_{4}\}$) are kept local. Router and expert parameters are updated iteratively using local validation and training datasets.
There has not yet been a solution to deal with both system heterogeneity and data heterogeneity. Towards this goal, we propose a novel Collaborative learning approach via a Mixture of Generalists and Specialists (CoMiGS). Our approach allows users to share part of the knowledge while keeping some knowledge user-specific, thus providing personalized solutions. We name the shared part generalists and the user-specific part specialists. Like all previous works, the generalists and specialists are simply LoRA modules. At the same time, as long as the shared part can be aggregated, the user-specific part can be of different sizes, which can be adapted to various device capacities, as illustrated by different numbers of specialists across users in Figure 1.
We integrate the expertise of generalists and specialists using a learned router that determines aggregation weights, following the Mixture-of-Experts (MoE) architecture (Fedus et al., 2022b). As in typical MoE designs for language modeling (Jiang et al., 2024; Fan et al., 2024), we also use tokens as the routing unit. Although users may have different topic preferences or linguistic styles, they still share common tokens in their vocabularies. Our goal is to route these shared tokens to the generalists so they can be jointly learned across users.
The closest work to ours is pFedMoE from Yi et al. (2024) from the vision domain, where each user has a shared homogeneous small feature extractor, a localized heterogeneous feature extractor, and a localized routing network, with routing unit being a semantic unit – an image. The three components are simultaneously updated. Compared to pFedMoE, our method CoMiGS introduces three key updates: 1) we reformulate the learning objective into a bi-level optimization framework, following the inherent hierarchical order of router and expert learning; 2) we refine the routing mechanism by using the smallest unit like a pixel in an image, which is a token; 3) unlike pFedMoE’s fixed two-expert limit per user, we support varying numbers of expert modules across users.
@@
  • We propose a novel approach (CoMiGS) for on-device personalized collaborative fine-tuning of LLMs, introducing an innovative bi-level formulation of the Mixture-of-Experts learning objective. Our approach can effectively tackle distribution shifts in local data.
@@
  • Our framework separates resource heterogeneity from data quantity. Users with larger local datasets benefit from a bigger model, while users with more powerful models but smaller datasets are less prone to overfitting.
  • We release a codebase [^1] Our code base is available at https://github.com/epfml/CoMiGS. for collaborative LLMs that allows users to easily define their own collaboration strategies, facilitating and advancing future research efforts in this field.
@@ 3 Method
We aim to improve personalized performance for each user on their target distributions, where distribution shifts can be allowed. Building on the hierarchical insights of MoE learning, we formulate our learning objective into a bi-level optimization problem, where expert parameters are learned using the relatively large-sized training sets, while routing parameters are updated using the small-sized validation sets. We further let experts diversify into generalists and specialists via parameter aggregation or localization, to leverage both collective power and specialized knowledge. As the problem solver, we provide a multi-round gradient-based algorithm.
@@ 3.1 Notions and problem setup
As illustrated in Figure 1, there are two sets of model parameters within each user: expert parameters, denoted as ${\mathbf{\Theta}}={\mathbf{\theta}}^{G}\cup\{{\mathbf{\theta}}_{i}^{S}\}$, where ${\mathbf{\theta}}^{G}$ is shared across the users and $\{{\mathbf{\theta}}_{i}^{S}\}$ are user-specific specialist parameters; and routing parameters, denoted as ${\mathbf{\Phi}}=\{{\mathbf{\phi}}_{i}\}$. $i\in\{1,2,..,N\}$ is the user index. Our ultimate goal is to optimize the average target performance across all users.
Our experts are simply LoRA modules, which approximate model updates $\Delta{\bm{W}}\in\mathbb{R}^{m\times n}$ with a multiplication of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{m\times r}$ and ${\bm{B}}\in\mathbb{R}^{r\times n}$ with rank $r\ll m,n$. ${\mathbf{\theta}}^{G}$ and ${\mathbf{\theta}}^{S}$ are disjoint sets of LoRA A and B matrices.
3.2 A bi-level formulation
Essentially, we adopt an MoE architecture, apart from aggregating certain expert parameters. Instead of learning routing and expert parameters simultaneously like the conventional way in LLMs (Zoph et al., 2022; Fedus et al., 2022a), we update the two sets of parameters in an alternating fashion. We observe a natural hierarchy between the experts and the router: the assignment of tokens to experts depends on the router’s outputs, while the experts’ parameters are updated based on the assigned tokens. In this way, the experts’ development follows the router’s decisions, establishing an inherent leader-follower structure. Following Von Stackelberg (2010), we formulate the hierarchical problem as a bi-level optimization objective in (1). Notably, one of the earliest MoE works (Jordan & Jacobs, 1994), also demonstrates a hierarchical structure, though for a probabilistic interpretation. In contrast, we approach the hierarchical structure from an optimization perspective, formulating the learning process as two nested optimization problems.
where ${\mathcal{L}}$ is the language modeling loss. Note we write ${\bm{X}}_{i}$ as the label here, as this is a self-supervised task. Labels are simply shifted inputs. The routing parameters ${\mathbf{\Phi}}=\{{\mathbf{\phi}}_{i}\}$ are updated based on the validation loss, which reflects the target distribution (outer optimization), while the expert parameters ${\mathbf{\Theta}}={\mathbf{\theta}}^{G}\cup\{{\mathbf{\theta}}^{S}_{i}\}$ are updated using the training loss (inner optimization). This formulation further brings in the following benefits: 1) routing parameters are smaller in size, making them easier to overfit. By separating the two losses, the routing parameters can be updated less frequently using the smaller validation set (a visual evidence of less frequent router update leading to improved performance is provided in Figure 7 in the Appendix); 2) this approach handles situations where target distributions differ from training distributions more effectively, as the router outputs (i.e., how the experts should be weighted) can be tailored to specific tasks.
@@ 3.2 A bi-level formulation
Input: Expert parameters $\{{\mathbf{\theta}}_{i,0}^{G},{\mathbf{\theta}}_{i,0}^{S}\}$, routing parameters ${\mathbf{\phi}}_{i,0}$}. Local training data and validation data $\{{\bm{X}}_{i}^{\text{train}},{\bm{X}}_{i}^{\text{valid}}\}$, $i\in\{1,2,..,N\}$. Communication round $T$ and routing update period $\tau$. Load balancing weight $\lambda$.
@@
     Each device $i\in\{1,2,..,N\}$ sends generalist weights ${\mathbf{\theta}}_{i,t}^{G}$ to the server
@@
3.3 Our Algorithm
To solve (1), we use a multi-round gradient-based algorithm as shown in Alg.1, where only generalist parameters are shared and aggregated, and specialist and router parameters are updated locally. While the scheme requires a server, it can alternatively be implemented in a serverless all2all fashion, which requires $N$ times more communication overhead and we do not further pursue this here.
@@ 3.3 Our Algorithm
Alternating update of two sets of parameters is a standard way to solve bi-level optimization problems (Chen et al., 2021). In between two communication rounds, we perform alternating updates of expert and routing parameters using local training and validation sets separately. The updates optimize the objectives given in (2) and (3) respectively. Since the updates of ${\mathbf{\Theta}}$ and ${\mathbf{\Phi}}$ are disentangled, they do not need to be updated at the same frequency. The routing parameters are smaller in size and thus can be updated less frequently.
When updating model parameters, we include an additional load-balancing term as in Fedus et al. (2022a), which is standard in MoE implementation and encourages even distribution of token assignments to experts. A discussion over the load balancing term is included in Appendix C.4. It is observed that a load-balancing term can improve test performance compared to not having one. However, directing more tokens to the generalists has no noticeable effect.
@@ 4.1.1 Datasets
We selected three diverse distributed datasets to demonstrate the efficacy of our proposed algorithm:
@@
    SlimPajama: We pick the following four categories – StackExchange, Github Codes, ArXiv, Book from Soboleva et al. (2023);
@@
Opting for the most challenging scenario, each user is assigned a unique category, as shown in Figure 8, where users can have varying data quantities. Given our emphasis on next token prediction, we anticipate shared predictions among users while maintaining category-specific distinctions. For details of our user data splits, please refer to Appendix B. We further create the following two scenarios to showcase the wide applicability of our method:
In-Distribution Tasks.
For each user, we construct validation and test datasets that follow the same distribution as the training data. We address two scenarios in this context: (i) variation in language usage across users (Multilingual Wikipedia), and (ii) variation in topic coverage across users (SlimPajama).
Out-of-Distribution Tasks.
For each user, we create validation and test datasets from a distribution different from the training data. During training, each user is assigned a single News category from AG News, but their validation and test sets consist of a uniform mixture of all categories. This approach accounts for potential shifts in topics within users.
@@ 4.1.2 Experimental Details
We start by equipping users with the same model architecture locally, to illustrate the effectiveness of our hierarchical learning of routing and expert parameters. We compare our one generalist one specialist (CoMiGS-1G1S) method to the following baselines. In order to match the trainable parameter count of our method, we use 2 times LoRA modules within each user.
@@
    • pFedMoE: We directly apply the method from Yi et al. (2024) in the language domain where we update routing and expert parameters at the same time and choose tokens as a routing unit.
@@
    Ablations:
    • CoMiGS-2S: Both of the LoRA experts are specialists, meaning their weights are neither shared nor aggregated. The routing parameters are updated using a separate validation set like in CoMiGS-1G1S.
    • CoMiGS-2G: Both of the LoRA experts are generalists, meaning their weights are always shared and aggregated. The routing parameters are updated using a separate validation set like in CoMiGS-1G1S.
In Distribution Out of Distribution
Multilingual SlimPajama AG News
Pretrained 156.12 37.19 90.65
Centralized 55.41 (0.12) 19.53 (0.14) 28.19 (0.52)
Local 54.38 (0.32) 26.95 (0.14) 41.46 (0.06)
FedAvg 58.80 (0.34) 23.27 (0.05) 31.84 (0.02)
PCL 54.53 (0.19) 26.99 (0.19) 32.25 (0.12)
pFedMoE 52.27 (0.17) 25.40 (0.09) 38.72 (0.21)
CoMiGS - 2S (ours) 46.36 (0.16) 22.51 (0.08) 35.81 (0.13)
CoMiGS - 2G (ours) 58.31 (0.17) 21.36 (0.01) 31.18 (0.05)
CoMiGS - 1G1S (ours) 47.19 (0.10) 21.79 (0.04) 33.53 (0.03)
Table 1: Mean test perplexity over the users with homogeneous models, averaged across 3 seeds. Mean (std) with a rank locator for the mean (the lower the better). Green denotes the best performing methods and red denotes our method. An extended version of the table can be found in Table 4.